# Crossed product splitting of intermediate operator algebras via $2$-cocycles

### v1

@@ 6. K0-obsrtuction: New tensor component $\mathcal{O}_{2}$ is necessary and minimal
We first see that the necessity of a new tensor component already appears in the following elementary and well-known example. Denote by $\mathbb{M}_{2^{\infty}}$ the CAR algebra. We identify $\mathbb{M}_{2^{\infty}}$ with the gauge fixed point algebra of $\mathcal{O}_{2}$ in the standard way. Then, as observed in [54], the inclusion $\mathbb{M}_{2^{\infty}}\subset\mathcal{O}_{2}$ is C-irreducible. Let $\alpha\colon\mathbb{Z}\curvearrowright\mathcal{O}_{2}$ be an automorphism commuting with the gauge action. (For instance, any quasi-free automorphism commutes with the gauge action.) Then for any $\chi\in\widehat{\mathbb{T}}\setminus\{1\}$, the set
@@
Next, by modifying the first construction, we also give examples which show that our adding tensor component $\mathcal{O}_{2}$ cannot be replaced by any other unital C-algebra $A$ with $[1_{A}]_{0}\neq 0$ in ${\rm K}_{0}(A)$. Before the construction, we state its consequence. Since any unital properly infinite C-algebra $A$ with $[1_{A}]_{0}=0$ contains $\mathcal{O}_{2}$ as a unital C-algebra (by [15]), this shows that our choice of the tensor component $\mathcal{O}_{2}$ is a minimal possible choice among regular unital simple C-algebras. We also note that, even considering pathological unital simple C-algebras (see [53]), the condition $[1_{B}]_{0}=0$ for a unital simple C-algebra $B$ forces the matrix amplification $\mathbb{M}_{k}(B)$ for a sufficiently large $k\in\mathbb{N}$ to contain $\mathcal{O}_{2}$ as a unital C-algebra. Thus there is no essential room of an improvement for the choice of $\mathcal{O}_{2}$.
@@
Let $s\in C$ denote the canonical implementing isometry element of $\rho$. Note that $ss^{\ast}=p$. Let $\sigma\colon\mathbb{T}\curvearrowright C$ be the gauge action. (Namely, the action given by $\sigma_{z}(b)=b$ and $\sigma_{z}(s)=zs$ for $b\in B$ and $z\in\mathbb{T}$.) We recall that $A$ is a unital C-algebra with $[1_{A}]_{0}\neq 0$. On $C$ and $A\otimes C$, we equip with the $\mathbb{T}$-actions $\sigma$ and $\mbox{\rm id}_{A}\otimes\sigma$ respectively. Let $\chi$ be the character on $\mathbb{T}$ given by $\chi_{z}=z$. Since one has the completely contractive projection ${\mathbf{E}}^{\mathbb{T}}_{\chi}\colon C\rightarrow C_{\chi}$, it follows from the description of the algebraic elements in the formula (*) in page 114 of [45] that $C_{\chi}=Bs$. This yields $(A\otimes C)_{\chi}=(A\otimes B)\cdot(1\otimes s)$. We show $(A\otimes C)_{\chi}\cap(A\otimes C)^{\rm u}=\emptyset$. Suppose that one has a unitary element $u$ of the form
@@
which sends $[1_{A}\otimes p]_{0}$ to $[1_{A}]_{0}\otimes\delta_{1}$ and $[1]_{0}$ to $[1_{A}]_{0}\otimes\delta_{0}$. Since $[1_{A}]_{0}\neq 0$, these two elements are different. This is a contradiction. Hence $(A\otimes C)_{\chi}\cap(A\otimes C)^{\rm u}=\emptyset$.
@@ 7. Von Neumann algebra case
Let $N\subset M$ be a von Neumann algebra inclusion. We say that $\alpha\in{\rm Aut}(M)$ with $\alpha(N)=N$ is centrally free for $N\subset M$ if there is a partition of unity $(p_{n})_{n\in\mathbb{N}}$ in $(N^{\omega}\cap M_{\omega})^{\rm p}$ with $p_{n}\alpha(p_{n})=0$ for $n\in\mathbb{N}$. We say that an inclusion $N\subset(M,\alpha,\mathfrak{w})$ of ((((twisted)))) $\Gamma$-von Neumann algebras is centrally $\Gamma$-free if every $\alpha_{s}$; $s\in\Gamma\setminus\{e\}$, is centrally free for $N\subset M$.
@@
Here for the last inequality, we use the following fact: For any $(q_{n})_{n\in\mathbb{N}}\in\ell^{\infty}(\mathbb{N},N)^{\rm p}$ representing an element of $N^{\omega}\cap M_{\omega}$, one has
@@
on $[0,\infty)$. Set $y:=1\otimes f(|d^{\ast}|)d\in\Theta({\mathbf{E}}_{s}(C))$. Then $y$ is a nonzero partial isometry element. Hence $yy^{\ast}$ and $y^{\ast}y$ are nonzero projections in $1_{\mathbb{B}}\otimes(C\cap M)_{p}$ and in $1_{\mathbb{B}}\otimes(\alpha_{s}(C\cap M))_{p}$ respectively. Since $(C\cap M)_{p}$ is a factor, by the Comparison Theorem, one can find isometry elements $v\in\mathbb{B}\mathbin{\bar{\otimes}}(C\cap M)_{p}$ and $w\in\mathbb{B}\mathbin{\bar{\otimes}}(\alpha_{s}(C\cap M))_{p}$ with $vv^{\ast}=yy^{\ast}$, $ww^{\ast}=y^{\ast}y$. These relations yield $\Theta^{-1}(v^{\ast}yw)\in{\mathbf{E}}_{s}(C)\cap M^{\rm u}$. ∎
@@
With Lemmas 7.3 to 7.6 in hand, now the proof is essentially the same as the unital C-algebra case. In the second case (the case $N$ is infinite), one can further eliminate the $2$-cocycle, by the stabilization trick [55]. ∎
@@
Similar to the C-algebra case (see Section 6), in general the statement of Theorem E fails. To see this, consider the state $\varphi:=\tau\circ{\mathbf{E}}^{\mathbb{T}}$ on $\mathcal{O}_{2}$. Here $\tau$ denotes the unique tracial state on $\mathbb{M}_{2^{\infty}}$. Let $N\subset M$ be the weak-$\ast$ closures of the C-algebras $\mathbb{M}_{2^{\infty}}\subset\mathcal{O}_{2}$ in the GNS representation of $\varphi$. Then $N$ is a type II1 factor, $M$ is a type III factor ([39]), and the inclusion is irreducible. As $\varphi$ is gauge invariant, the gauge action extends to the action $\mathbb{T}\curvearrowright M$ with $M^{\mathbb{T}}=N$. Let $\alpha\colon\mathbb{Z}\curvearrowright M$ be any action commuting with the gauge action. (For instance, any quasi-free action on $\mathcal{O}_{2}$ extends to such an action.) Fix $\chi\in\widehat{\mathbb{T}}\setminus\{1\}$. We set
@@ 8. Galois’s type result for compact-by-discrete groups
For a C-algebra $A$, let $A_{\infty}:=\ell^{\infty}(\mathbb{N},A)/c_{0}(\mathbb{N},A)$. For a continuous group action $\alpha\colon G\curvearrowright A$, let $\alpha_{\infty}\colon G\curvearrowright A_{\infty}$ denote the induced (possibly discontinuous) action. Denote by $A_{\infty,\alpha}$ the set of all $G$-continuous elements with respect to $\alpha_{\infty}$.
@@
Let $\alpha\colon K\rtimes_{\sigma}\Gamma\curvearrowright A$ be an isometrically shift-absorbing action on a separable C-algebra $A$. Then $A^{K}\subset(A,\alpha|_{\Gamma})$ is centrally $\Gamma$-free.
@@
By a characterization of the isometrically shift-absorption ([18], Proposition 3.8 (iv)), we have a ($K\rtimes_{\sigma}\Gamma$)-equivariant $A$-bilinear map
@@
Thanks to Lemmas 8.2 and 8.3 and the Galois correspondence for $A^{K}\subset A$, clearly the map in the statement is injective. The relation $(L_{1},\Lambda_{1},\omega_{1})\preceq(L_{2},\Lambda_{2},\omega_{2})$ for two elements in $\mathcal{L}(K,\Gamma,\sigma)$ implies $\Lambda_{1}<\Lambda_{2}$ and $A_{\omega_{1}(s)}\subset A_{\omega_{2}(s)}$ for all $s\in\Lambda_{1}$. Hence $\operatorname{{\rm C}^{\ast}}(L_{1},\Lambda_{1},\omega_{1})\subset\operatorname{{\rm C}^{\ast}}(L_{2},\Lambda_{2},\omega_{2})$. Conversely, if $\operatorname{{\rm C}^{\ast}}(L_{1},\Lambda_{1},\omega_{1})\subset\operatorname{{\rm C}^{\ast}}(L_{2},\Lambda_{2},\omega_{2})$, one clearly has $A^{L_{1}}\subset A^{L_{2}}$ and $\Lambda_{1}\subset\Lambda_{2}$ (by Lemma 8.3). By the Galois correspoindence [35], the first relation implies $L_{2}<L_{1}$. Also, for any $s\in\Lambda_{1}$, one has

### v2

@@ 6. K0-obstruction: New tensor component $\mathcal{O}_{2}$ is necessary and minimal
[…]
@@
Next, by modifying the first construction, we also give examples which show that our adding tensor component $\mathcal{O}_{2}$ cannot be replaced by any other unital C-algebra AA with $1_{A}]_{0}\neq 0$ in ${\rm K}_{0}(A)$. Before the construction, we state its consequence. Since any unital properly infinite C-algebra AA with $1_{A}]_{0}=0$ contains $\mathcal{O}_{2}$ as a unital C-algebra (by [15]), this shows that our choice of the tensor component $\mathcal{O}_{2}$ is a minimal possible choice among regular unital simple C-algebras. We also note that, even considering pathological unital simple C-algebras (see [53]), the condition $1_{B}]_{0}=0$ for a unital simple C-algebra $B$ forces the matrix amplification $\mathbb{M}_{k}(B)$ for a sufficiently large $k\in\mathbb{N}$ to contain $\mathcal{O}_{2}$ as a unital C-algebra. Thus there is no essential room of an improvement for the choice of $\mathcal{O}_{2}$.
@@
Let $s\in C$ denote the canonical implementing isometry element of $\rho$. Note that $ss^{\ast}=p$. Let $\sigma\colon\mathbb{T}\curvearrowright C$ be the gauge action. (Namely, the action given by $\sigma_{z}(b)=b$ and $\sigma_{z}(s)=zs$ for $b\in B$ and $z\in\mathbb{T}$.) We recall that $A$ is a unital C-algebra with $1_{A}]_{0}\neq 0$. On $C$ and $A\otimes C$, we equip with the $\mathbb{T}$-actions $\sigma$ and $\mbox{\rm id}_{A}\otimes\sigma$ respectively. Let $\chi$ be the character on $\mathbb{T}$ given by $\chi_{z}=z$. Since one has the completely contractive projection ${\mathbf{E}}^{\mathbb{T}}_{\chi}\colon C\rightarrow C_{\chi}$, it follows from the description of the algebraic elements in the formula (*) in page 114 of [45] that $C_{\chi}=Bs$. This yields $(A\otimes C)_{\chi}=(A\otimes B)\cdot(1\otimes s)$. We show $(A\otimes C)_{\chi}\cap(A\otimes C)^{\rm u}=\emptyset$. Suppose that one has a unitary element $u$ of the form
@@
which sends $1_{A}\otimes p]_{0}$ to $1_{A}]_{0}\otimes\delta_{1}$ and $1]_{0}$ to $1_{A}]_{0}\otimes\delta_{0}$. Since $1_{A}]_{0}\neq 0$, these two elements are different. This is a contradiction. Hence $(A\otimes C)_{\chi}\cap(A\otimes C)^{\rm u}=\emptyset$.
@@ 7. Von Neumann algebra case
Let $N\subset M$ be a von Neumann algebra inclusion. We say that $\alpha\in{\rm Aut}(M)$ with $\alpha(N)=N$ is centrally free for $N\subset M$ if there is a partition of unity $(p_{n})_{n\in\mathbb{N}}$ in $(N^{\omega}\cap M_{\omega})^{\rm p}$ with $p_{n}\alpha(p_{n})=0$ for $n\in\mathbb{N}$. We say that an inclusion $N\subset(M,\alpha,\mathfrak{w})$ of ((twisted)) $\Gamma$-von Neumann algebras is centrally $\Gamma$-free if every $\alpha_{s}$; $s\in\Gamma\setminus\{e\}$, is centrally free for $N\subset M$.
@@
Here for the last inequality, we use the following fact: For any $(q_{n})_{n\in\mathbb{N}}\in\ell^{\infty}(\mathbb{N},N)^{\rm p}$ representing an element of $N^{\omega}\cap M_{\omega}$ and any $a\in M\mathop{{\bar{\rtimes}}_{\alpha,\mathfrak{w}}}\Gamma$, one has
@@
on $0,\infty)$. Set $y:=1\otimes f(|d^{\ast}|)d\in\Theta({\mathbf{E}}_{s}(C))$. Then $y$ is a nonzero partial isometry element. Hence $yy^{\ast}$ and $y^{\ast}y$ are nonzero projections in $1_{\mathbb{B}}\otimes(C\cap M)_{p}$ and in $1_{\mathbb{B}}\otimes(\alpha_{s}(C\cap M))_{p}$ respectively. Since $(C\cap M)_{p}$ is a factor, by the Comparison Theorem, one can find isometry elements $v\in\mathbb{B}\mathbin{\bar{\otimes}}(C\cap M)_{p}$ and $w\in\mathbb{B}\mathbin{\bar{\otimes}}(\alpha_{s}(C\cap M))_{p}$ with $vv^{\ast}=yy^{\ast}$, $ww^{\ast}=y^{\ast}y$. These relations yield $\Theta^{-1}(v^{\ast}yw)\in{\mathbf{E}}_{s}(C)\cap M^{\rm u}$. ∎
@@
With Lemmas 7.3 to 7.6 in hand, now the proof is essentially the same as the unital C-algebra case. In the second case (the case NN is infinite), one can further eliminate the 22-cocycle, by the stabilization trick [55]. ∎
@@
[…]
@@ 8. Galois’s type result for compact-by-discrete groups
For a C-algebra AA, let $A_{\infty}:=\ell^{\infty}(\mathbb{N},A)/c_{0}(\mathbb{N},A)$. For a continuous group action $\alpha\colon G\curvearrowright A$, let $\alpha_{\infty}\colon G\curvearrowright A_{\infty}$ denote the induced (possibly discontinuous) action. Denote by $A_{\infty,\alpha}$ the set of all $G$-continuous elements with respect to $\alpha_{\infty}$.
@@
Let $\alpha\colon K\rtimes_{\sigma}\Gamma\curvearrowright A$ be an isometrically shift-absorbing action on a separable C-algebra AA. Then $A^{K}\subset(A,\alpha|_{\Gamma})$ is centrally $\Gamma$-free.
@@
By a characterization of the isometrically shift-absorption ([18], Proposition 3.8 (iv)), we have a ($K\rtimes_{\sigma}\Gamma$)-equivariant AA-bilinear map
@@
Thanks to Lemmas 8.2 and 8.3 and the Galois correspondence for $A^{K}\subset A$, clearly the map in the statement is injective. The relation $(L_{1},\Lambda_{1},\omega_{1})\preceq(L_{2},\Lambda_{2},\omega_{2})$ for two elements in $\mathcal{L}(K,\Gamma,\sigma)$ implies $\Lambda_{1}<\Lambda_{2}$ and $A_{\omega_{1}(s)}\subset A_{\omega_{2}(s)}$ for all $s\in\Lambda_{1}$. Hence $\operatorname{{\rm C}^{\ast}}(L_{1},\Lambda_{1},\omega_{1})\subset\operatorname{{\rm C}^{\ast}}(L_{2},\Lambda_{2},\omega_{2})$. Conversely, if $\operatorname{{\rm C}^{\ast}}(L_{1},\Lambda_{1},\omega_{1})\subset\operatorname{{\rm C}^{\ast}}(L_{2},\Lambda_{2},\omega_{2})$, one clearly has $A^{L_{1}}\subset A^{L_{2}}$ and $\Lambda_{1}\subset\Lambda_{2}$ (by Lemma 8.3). By the Galois correspondence [35], the first relation implies $L_{2}<L_{1}$. Also, for any $s\in\Lambda_{1}$, one has
@@ Acknowledgements
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.